# HARMONIOUS – Human-like reactive motion control and multimodal perception for humanoid robots

### v1

@@ Abstract
For safe and effective operation of humanoid robots in human-populated environments, the problem of commanding a large number of Degrees of Freedom (DoF) while simultaneously considering dynamic obstacles and human proximity has still not been solved. We present a new reactive motion controller that commands two arms of a humanoid robot and three torso joints (17 DoF in total). We formulate a quadratic program that seeks joint velocity commands respecting multiple constraints while minimizing the magnitude of the velocities. We introduce a new unified treatment of obstacles that dynamically maps visual and proximity (pre-collision) and tactile (post-collision) obstacles as additional constraints to the motion controller, in a distributed fashion over surface of the upper-body of the iCub robot (with 2000 pressure-sensitive receptors). The bio-inspired controller: (i) produces human-like minimum jerk movement profiles; (ii) gives rise to a robot with whole-body visuo-tactile awareness, resembling peripersonal space representations. The controller was extensively experimentally validated, including a physical human-robot interaction scenario.
@@ I Introduction
We present a reactive motion controller that commands the upper body of the iCub humanoid robot (2x 7 Degrees of Freedom (DoF) in every arm; 3 torso DoF), whereby the two hands can have separate tasks in Cartesian space or a common task (see Fig.1). Unique to our approach, obstacles perceived in three different modalities—visual and proximity (pre-collision) and tactile (post-collision)—are dynamically aggregated and projected first onto locations on individual body parts and then remapped as constraints into the joint velocity space.
@@
I-A Inverse kinematics
With robot tasks are defined in Cartesian (also called operational or task) space while they are commanded in joint space, some form of inverse kinematics (IK) mapping ($\mathbf{q}=f(\mathbf{x})$) is indispensable. For industrial manipulators, a closed-form analytical solution is often available, especially for 6 DoF arms with a so-called spherical wrist. Closed-form solutions are global and fast to compute. However, such analytical solutions exploit specific geometric relationships and are in general not available for redundant robots with an arbitrary kinematic structure (but cf. [1]). Numerical approaches to IK come with a higher computational load, but are more general—applicable to any kinematic chain. TRAC-IK [2] is a state-of-the-art IK solver that solves IK as a sequential quadratic programming problem. Lloyd et al. [3] presented an IK algorithm that uses the third-order root finding method to increase the speed and robustness of finding IK solutions. The solution consists of a set of target joint angles only—smooth trajectory generation and execution has to be taken care of separately. Pattacini et al. [4] combined a trajectory generator, a non-linear IK solver, and a Cartesian controller to ensure human-like minimum jerk movements in both Cartesian and joint space. None of the above methods can handle dynamic obstacles in the workspace. An overview of the related work from this and the following sections is in Tab. I.
Work Control problem ROA
@@ I-A Inverse kinematics
NPR OEM ETS
Ours DiffK - QP problem LIC 3 - vision, proximity, touch 17 Velocity damping Yes Yes Yes
@@
[5] DiffK - Damped LS LIC 0 - known obstacles 8 Damped LS No Yes No
[6] DiffK - Pseudoinverse PF 0 - known obstacles 7 No No Yes Yes
[7] DiffK - Pseudoinverse RV 1 - vision 7 No No No No
[8] DiffK - QP problem LIC 0 - known obstacles 7 Maximize manipulability No Yes No
[9] DiffK - QP problem LIC 2 - proximity, touch 7 Velocity damping Yes No No
[10] DiffK - QP problem LIC 1 - proximity 7 Velocity damping No No No
[11] DiffK - QP problem LIC 0 - known obstacles 7 No No Yes No
[12] DiffK - QP problem - - 12 Maximize manipulability No Yes No
[13] DiffK - QP problem LIC 0 - known obstacles 6 No No Yes No
[14] IK - Nonlinear opt. CF 0 - known obstacles 8 Maximize condition number No Yes No
[15] Operational space control PF 0 - known obstacles 6 No No Yes No
[16] ERG + PD control PF 1 - vision 7 No No Yes No
[17] IK - Nonlinear opt. RV 2 - vision, touch 10 No No Yes Yes
TABLE I: Comparison of different approaches to solving the IK problem. Control problem acronyms are: differential kinematics (DiffK), inverse kinematics (IK), quadratic programming (QP), least squares (LS), and explicit reference governor (ERG). Reactive obstacle avoidance (ROA) types are linear inequality constraint (LIC), self-motion (SM), potential field (PF), danger index (DI), repulsive vector (RV), and part of the cost function (CF). The last columns are natural position reward (NPR), orientation error minimization (OEM), and explicit trajectory sampling (ETS).
I-B Obstacle avoidance and differential kinematics
@@
Alternatively, the motion control task can be defined using the forward differential formulation, searching for joint velocities $\mathbf{\dot{q}}$ that best match desired Cartesian velocity of the end effector $\mathbf{\dot{x}}$, thus minimizing $||\mathbf{\dot{x}}-\mathbf{J}(\mathbf{q})\mathbf{\dot{q}}||$. To this end, it is possible to use quadratic programming (QP) optimization with a quadratic cost function and linear constraints [8, 9, 10, 11, 12, 13, 26]. To avoid singular configurations, two options are usually given: add manipulability maximization to the cost function [8, 12] or dampen joint velocities when close to singularities [9, 10, 27]. In addition, undesirable configurations can be avoided by extending the optimization cost function to motivate the robot to be close to the middle of the joint limits [9] or close to the robot resting position [4]. In the case of the QP problem, obstacles are represented as linear inequality constraints allowing motions in tangential directions improving target reachability [8, 9, 10].
Most of the mentioned works do not solve self-collisions explicitly. The exceptions are [10], where self-collisions are solved by adding additional obstacle points for robot links and the end effector, and [14], where a neural network is trained to favor configurations far from self-collision states. Unlike most of the mentioned controllers that presented solutions of the IK problem for a single-arm robot, Tong et al. [12] proposed a four-criterion-optimization coordination motion scheme for a dual-arm robot that deals with coordination constraints and physical constraints. However, their approach does not consider obstacle avoidance and the torso joints of the dual-arm robot.
@@
I-C Human-like motion generation
Several controllers can work as explicit local trajectory generators for sampling the trajectory between the start and target pose. The algorithm in [29] combines control with path planning and can handle arbitrary desired velocity profiles for the robot. The Cartesian controller [4] produces human-like quasi-straight trajectories whose typical characteristic is to rely on minimum-jerk profiles. A reactive controller in [17, 30] uses the minimum-jerk trajectory sampling as in [4]. Suleiman [13] proposed an algorithm for implicit minimum jerk trajectories by replacing joint velocities by joint jerk as control parameters in the optimization problem.
I-D Sensing of obstacles and contacts
@@
To sense at a distance, RGB-D cameras are frequently adopted [7, 31, 32, 33, 34]. Depth information can be used to compute a set of spheres that represent obstacles [33] or to create a depth space representation that is used to generate repulsive vectors for the robot [7]. Magrini and De Luca [24] estimated the contact force based on the contact point detected by a Kinect sensor. Alternative visual sensing for obstacle detection are tracking systems that follow markers attached to specific body parts, e.g., a wrist [29] or arm joints and a head [16]. Aljalbout et al. [35] used a static RGB camera in simulation and trained a convolutional neural network for obstacle avoidance.
@@
Collisions with a moving robot may be acceptable provided that the contact forces are within limits. Collisions need to be detected, located, identified (e.g., determining the forces), and, possibly following classification of the collision, reacted upon (e.g., stop, slow down, retract, etc.) [36]. Force/torque sensors within the robot structure, together with models of its dynamics, can be employed. Alternatively, contacts can be perceived through tactile sensors. We will focus on large area sensitive skins. Differently than when force/torque sensors are employed, collision localization/isolation is easier, even for multi-contacts. Cirillo et al. [23] developed a flexible skin, allowing measurement of contact position and three components of contact force. Calandra et al. [37] presented a data-driven method for fast and accurate prediction of joint torques from contacts detected by tactile skin. Albini et al. used robot skin feedback HRI [38] and robot control in a cluttered environment [22]. Kuehn and Haddadin [39] introduced an artificial Robot Nervous System, which integrates tactile sensation and reflex reactions into robot control through the concept of robot pain sensation. The pain-reflex movements, in combination with the new skins, can improve the human feeling of robots and their safety. Some large area sensitive skins have been safety-rated and are deployed in collaborative robot applications (see Airskin and the analysis of its effects in [40]).
@@
Individual visual sensors—on the robot itself or in other locations in the workspace—are prone to occlusions (which can be mitigated by employing multiple sensors [41]). Large area tactile arrays are distributed over the whole robot body, but only respond after physical contact. Distributed proximity sensors provide an alternative—sensing at a small distance over the whole robot body. Proximity [9, 10] or tomographic [42] sensors can be used for collision anticipation and contactless motion guidance.
@@
Typically, only one sensor type is used and the corresponding robot control is more or less dependent on the nature of the sensory information. Works combining distal and contact sensing, like vision and touch [31, 43] or proximity and touch [9] are scarce. Humans posses a dynamically established protective safety margin around the whole body, drawing on visuo-tactile interactions: Peripersonal space (PPS), e.g., [44]. Roncone et al. [43, 45] took inspiration in the PPS and implemented visuo-tactile receptive fields on a humanoid robot. Nguyen et al. [17, 30] deployed this in HRI to feed a reactive motion controller. As affordable and increasingly accurate sensors of different types are becoming ubiquitous, it is important and timely to develop a unified treatment of their signals such that they can be processed by a single motion controller. In this work we—to our knowledge for the first time—present a framework in which RGB-D, proximity, and tactile sensors are simultaneously processed, related to the complete robot body, and used to generate real-time constraints for a whole-body reactive motion controller.
I-E Proposed solution (HARMONIOUS) and contribution over state of the art
Here we present HARMONIOUS – Human-like reactive motion control and multimodal perception for humanoid robots. An overview is provided in Fig. 2. The core of the solution is the motion controller (red box) and the actual QP formulation (magenta) which is online fed by the target trajectory generator (blue box) and the constraints projected from obstacles (green box). A detailed description is provided in the Materials and Methods section. To the extent that this is possible, the main components of our solution are categorized and contrasted with representative works from the literature in Tab. I.
@@ I-E Proposed solution (HARMONIOUS) and contribution over state of the art
We formulate the control problem (second column in Tab. I) as a quadratic program in velocity space— employing differential kinematics, similarly to many recent works in this area [8, 9, 10, 11, 12, 13]. Reactive obstacle avoidance (ROA, second column) is implemented using linear inequality constraint (LIC). Importantly, this work is to our knowledge the first to employ three different sensory modalities (vision, proximity, touch), in a unified manner, to perceive and avoid obstacles (fourth column). The fifth column (DoF shown) lists the number of DoF employed in our approach and in the literature. HARMONIOUS handles kinematic singularities through velocity damping. The robot is rewarded for assuming a natural posture (NPR), unlike most of the works in the literature. HARMONIOUS can handle position and orientation targets (OEM) and samples the trajectory of end-effector positions (ETS), inspired by human reaching movements—quasi-straight trajectories with bell-shaped velocity profiles. On the other hand, the end-effector orientation sampling is calculated using a computer graphics method slerp for constant angular velocities, ensuring smooth rotations and the shortest path between initial and final orientations. In summary, HARMONIOUS employs state-of-the-art control problem and reactive obstacle avoidance formulation, but moves beyond the state of the art in providing a unified treatment of three different sensory modalities, the number of DoF controlled, and incorporation of natural position reward and orientation error minimization. Beyond what could fit in Tab. I, HARMONIOUS also features self-collision avoidance and the possibility of setting targets (tasks) for the two robot arms independently or jointly (bimanual task). To our knowledge, HARMONIOUS is unparalleled in the scale of the problem and the number of features it can tackle.
HARMONIOUS was tested against the most relevant solutions from related works [4, 8, 17] in simulation experiments without and with obstacles and achieved the best results. Then, it was verified with a real robot in several collision avoidance scenarios. Another experiment showed the successful cooperation of both arms in a safe bimanual task. Moreover, it was shown that the problem formulation can be easily extended with other task constraints. As the final evaluation, we tested HARMONIOUS in an interactive board game scenario with the human player dynamically perceived, simultaneously processing keypoints on its body, proximity signals, and physical contacts on the whole robot body.
II Materials and Methods
II-A Experimental setup

### v2

@@ Abstract
For safe and effective operation of humanoid robots in human-populated environments, the problem of commanding a large number of Degrees of Freedom (DoF) while simultaneously considering dynamic obstacles and human proximity has still not been solved. We present a new reactive motion controller that commands two arms of a humanoid robot and three torso joints (17 DoF in total). We formulate a quadratic program that seeks joint velocity commands respecting multiple constraints while minimizing the magnitude of the velocities. We introduce a new unified treatment of obstacles that dynamically maps visual and proximity (pre-collision) and tactile (post-collision) obstacles as additional constraints to the motion controller, in a distributed fashion over the surface of the upper body of the iCub robot (with 2000 pressure-sensitive receptors). This results in a bio-inspired controller that: (i) gives rise to a robot with whole-body visuo-tactile awareness, resembling peripersonal space representations, and (ii) produces human-like minimum jerk movement profiles. The controller was extensively experimentally validated, including a physical human-robot interaction scenario.
Index Terms:
Motion control, collision avoidance, humanoid robots, human-robot interaction
@@ I Introduction
We present a reactive motion controller that commands the upper body of the iCub humanoid robot ($2\times 7$ Degrees of Freedom (DoFs) in every arm; $3$ torso DoFs), whereby the two hands can have separate tasks in Cartesian space or a common task (see Fig.1). Unique to our approach, obstacles perceived in three different modalities—visual and proximity (pre-collision) and tactile (post-collision)—are dynamically aggregated and projected first onto locations on individual body parts and then remapped as constraints into the joint velocity space.
@@
Contributions. We have developed HARMONIOUS, a real-time reactive motion control system for upper body control of a humanoid robot sensing and avoiding contact with humans in its close proximity. Our specific contributions are the following.
  1. 1.
    HARMONIOUS is human-like in two important aspects: (i) it employs multimodal sensing around the whole body of the robot, “visuo-tactile awareness” resembling peripersonal space representations; (ii) it produces minimum jerk movement profiles which are characteristic of human motion. Together, these provide a basis for safe and natural human-robot interaction (HRI)—physical and social.
  2. 2.
    We developed a unified representation of the space around the robot that feeds a reactive motion controller. Dynamically moving obstacles ranging from physical contact with the robot (zero distance, perceived through touch) to obstacles in close or far proximity can be remapped onto the robot’s body parts, weighted, and transformed into kinematic constraints. Thus, there is a dense and multimodal representation of a peripersonal space analogue, which we refer to as the “perirobot space” (see [1] for more details). It is demonstrated here by combining touch, proximity and visual sensors, but other contact-related or range-based sensors could be used.
  3. 3.
    We have extensively and experimentally evaluated HARMONIOUS and demonstrated its superior performance to state-of-the-art robot controllers where possible ([2, 3, 4]). Moreover, to our knowledge, this is the first work to show real-time control of a humanoid robot upper torso (in total $17$ DoFs, with the possibility of different tasks for the two arms) faced with tens of dynamically moving obstacles around the whole robot body surface. Furthermore, kinematic singularities are handled through velocity damping and preferred postures are rewarded in the optimization problem formulation (see Tab. I for an overview). As the final evaluation, we tested HARMONIOUS in an interactive board game scenario with the human player dynamically perceived, simultaneously processing keypoints on his body, proximity signals, and physical contacts on the whole robot body.
  4. 4.
    The problem formulation is highly modular and both the minimization criteria (e.g., motivating preferred postures) or constraints to the quadratic program (e.g., obstacles) can be easily removed or added on the run.
II Related work
II-A Inverse kinematics
Robot tasks are defined in Cartesian (also called operational or task) space while they are commanded in joint space, some form of inverse kinematics (IK) mapping ($\mathbf{q}=f(\mathbf{x})$) is indispensable. For industrial manipulators, a closed-form analytical solution is often available, especially for 6 DoF arms with a so-called spherical wrist. Closed-form solutions are global and fast to compute. However, such analytical solutions exploit specific geometric relationships and are generally not available for redundant robots with an arbitrary kinematic structure ([5] is an exception). Numerical approaches to IK come with a higher computational load, but are more general and applicable to any kinematic chain. TRAC-IK [6] is a state-of-the-art IK solver that solves IK as a sequential quadratic programming problem. Lloyd et al. [7] presented an IK algorithm that uses the third-order root finding method to increase the speed and robustness of finding IK solutions. The solution consists of a set of target joint angles only—smooth trajectory generation and execution have to be taken care of separately. Pattacini et al. [8] combined a trajectory generator, a non-linear IK solver, and a Cartesian controller to ensure human-like minimum jerk movements in both Cartesian and joint space. None of the above methods can handle dynamic obstacles in the workspace. An overview of the related work from this and the following sections is in Tab. I.
Work Control problem
React. obs.
avoidance
@@ II-A Inverse kinematics
Pref. posture
motivation
Ori. error
minim.
Explicit
traj samp.
Ours DiffK - QP problem Lin. ineq. constr 3 - vision, proximity, touch 17 Velocity damping Yes Yes Yes
@@
[9] DiffK - Damped LS Lin. ineq. constr. 0 - known obstacles 8 Damped LS No Yes No
[10] DiffK - Pseudoinverse Potential field 0 - known obstacles 7 No No Yes Yes
[11] DiffK - Pseudoinverse Repulsive vector 1 - vision 7 No No No No
[2] DiffK - QP problem Lin. ineq. constr. 0 - known obstacles 7 Max. manipulability No Yes No
[12] DiffK - QP problem Lin. ineq. constr. 2 - proximity, touch 7 Velocity damping Yes No No
[13] DiffK - QP problem Lin. ineq. constr. 1 - proximity 7 Velocity damping No No No
[14] DiffK - QP problem Lin. ineq. constr. 0 - known obstacles 7 No No Yes No
[15] DiffK - QP problem - - 12 Max. manipulability No Yes No
[16] DiffK - QP problem Lin. ineq. constr. 0 - known obstacles 6 No No Yes No
[17] IK - Nonlinear opt. Cost function 0 - known obstacles 8 Max. condition number No Yes No
[18] Operational space control Potential field 0 - known obstacles 6 No No Yes No
[19] ERG + PD control Potential field 1 - vision 7 No No Yes No
[3] IK - Nonlinear opt. Repulsive vector 2 - vision, touch 10 No No Yes Yes
TABLE I: Comparison of different approaches to solving the inverse kinematics problem. Control problem acronyms are: differential kinematics (DiffK), inverse kinematics (IK), quadratic programming (QP), least squares (LS), and explicit reference governor (ERG).
II-B Obstacle avoidance and differential kinematics
@@
Alternatively, the motion control task can be defined using the forward differential formulation, searching for joint velocities $\mathbf{\dot{q}}$ that best match the desired Cartesian velocity of the end effector $\mathbf{\dot{x}}$, thus minimizing $||\mathbf{\dot{x}}-\mathbf{J}(\mathbf{q})\mathbf{\dot{q}}||$. To this end, it is possible to use quadratic programming (QP) optimization with a quadratic cost function and linear constraints [2, 12, 13, 14, 15, 16, 28]. To avoid singular configurations, two options are usually given: add manipulability maximization to the cost function [2, 15] or dampen joint velocities when close to singularities [12, 13, 29]. In addition, undesirable configurations can be avoided by extending the optimization cost function to motivate the robot to be close to the middle of the joint limits [12] or close to the robot resting position [8]. In the case of the QP problem, obstacles are represented as linear inequality constraints allowing motions in tangential directions improving target reachability [2, 12, 13].
Most of the mentioned works do not handle self-collisions explicitly. The exceptions are [13], where self-collisions are solved by adding additional obstacle points for robot links and the end effector, and [17], where a neural network is trained to favor configurations far from self-collision states. Moreover, De Santis et al. [30] presented an algorithm for real-time generation of self-collsion avoidance motions using repulsion forces converted to avoidance torques. Unlike most of the mentioned controllers that presented solutions of the IK problem for a single-arm robot, Tong et al. [15] proposed a four-criterion-optimization coordination motion scheme for a dual-arm robot that deals with coordination constraints and physical constraints. However, their approach does not consider obstacle avoidance and the torso joints of the dual-arm robot.
@@
II-C Human-like motion generation
Several controllers can work as explicit local trajectory generators to sample the trajectory between the start and the target pose. The algorithm in [32] combines control with path planning and can handle arbitrary desired velocity profiles for the robot. The Cartesian controller [8] produces human-like quasi-straight trajectories with minimum-jerk profiles characteristic of human motions. A reactive controller in [3, 4] uses the minimum-jerk trajectory sampling as in [8]. Suleiman [16] proposed an algorithm for implicit minimum jerk trajectories by replacing joint velocities by joint jerk as control parameters in the optimization problem.
II-D Sensing of obstacles and contacts
@@
To sense at a distance, RGB-D cameras are frequently adopted [11, 33, 34, 35, 36]. Depth information can be used to compute a set of spheres that represent obstacles [35] or to create a depth space representation that is used to generate repulsive vectors for the robot [11]. Magrini and De Luca [26] estimated the contact force based on the contact point detected by a Kinect sensor. Alternative visual sensing for obstacle detection are tracking systems that follow markers attached to specific body parts, e.g., wrist [32] or arm joints and head [19]. Aljalbout et al. [37] used a static RGB camera in simulation and trained a convolutional neural network for obstacle avoidance.
@@
Collisions with a moving robot may be acceptable provided that the contact forces are within limits. Collisions need to be detected, located, identified (e.g., determining the forces), and, possibly following classification of the collision, reacted upon (e.g., stop, slow down, retract, etc.) [38]. Force/torque sensors within the robot structure, together with models of its dynamics, can be employed. Alternatively, contacts can be perceived through tactile sensors. We will focus on large-area sensitive skins. Unlike when force/torque sensors are employed, collision localization/isolation is easier, even for multi-contacts. Cirillo et al. [25] developed a flexible skin, allowing measurement of contact position and three components of contact force. Calandra et al. [39] presented a data-driven method for a fast and accurate prediction of joint torques from contacts detected by tactile skin. Albini et al. used robot skin feedback for HRI [40] as well as robot control in a cluttered environment [24]. Kuehn and Haddadin [41] introduced an artificial robot nervous system, which integrates tactile sensation and reflex reactions into robot control through the concept of robot pain sensation. Some large-area sensitive skins have been safety-rated and are deployed in collaborative robot applications (see Airskin and the analysis of its effects in [42]).
@@
Individual visual sensors—on the robot itself or in other locations in the workspace—are prone to occlusions (which can be mitigated by employing multiple sensors [43]). Large area tactile arrays are distributed over the whole robot body, but only respond after physical contact. Distributed proximity sensors provide an alternative: sensing at a small distance over the whole robot body. Proximity [12, 13] or tomographic [44] sensors can be used for anticipation of collisions and contactless motion guidance.
Typically, only one sensor type is used, and the corresponding robot control is more or less dependent on the nature of the sensory information. On the other hand, there is a field of multimodal sensor fusion. Dean-Leon et al. [45] presented a fusion of pressure and proximity data from multimodal artificial skin to change the robot’s dynamic behavior, e.g., making a stiff robotic system compliant. Hierarchical control is often used together with the integration of several sensors, such as an artificial skin for collision avoidance with a camera for position-based visual servoing [46], or a LIDAR for collision avoidance and a multimodal skin for tactile interaction or distance/force control [47]. In addition to collision avoidance, fusion can be used for other tasks, such as opening the door [48]. Multimodal fusion can integrate more than sensor data as in [49], where contact location is estimated using a contact hypotheses fusion approach from multiple sensor modalities, the context of interaction, and the environment.
@@
Humans posses a dynamically established protective safety margin around the whole body, drawing on visuo-tactile interactions: Peripersonal space (PPS), e.g., [50]. Roncone et al. [51, 52] took inspiration from PPS and implemented visuo-tactile receptive fields on a humanoid robot. Nguyen et al. [3, 4] deployed this in HRI to feed a reactive motion controller. As affordable and increasingly accurate sensors of different types are becoming ubiquitous, it is important and timely to develop a unified treatment of their signals such that they can be processed by a single motion controller (see also [1] where a unified treatment of the perirobot space for collaborative robot applications is proposed).
II-E Comparison of HARMONIOUS with related works
In this work, we present a framework in which RGB-D, proximity, and tactile sensors are simultaneously processed, related to the complete robot body, and used to generate real-time constraints for a whole-body reactive motion controller.
To the extent that this is possible, the main components of our solution are categorized and contrasted with representative works from the literature in Tab. I.
@@ II-E Comparison of HARMONIOUS with related works
We formulate the control problem (second column in Tab. I) as a quadratic program in velocity space—employing differential kinematics, similarly to many recent works in this area [2, 12, 13, 14, 15, 16]. Reactive obstacle avoidance (second column) is implemented using a linear inequality constraint (LIC). Importantly, this work is to our knowledge the first to employ three different sensory modalities (vision, proximity, touch), in a unified manner, to perceive and avoid obstacles (fourth column). The fifth column (DoF shown) lists the number of DoF employed in our approach and in the literature. HARMONIOUS handles kinematic singularities through velocity damping. The robot is motivated to favor joint positions near a preferred posture, e.g., the center of the joint ranges or a home configuration to prevent undesirable configurations, unlike most of the works in the literature. HARMONIOUS can handle position and orientation targets (Orientation error minimization) and samples the trajectory of end-effector positions (Explicit trajectory sampling), inspired by human reaching movements—quasi-straight trajectories with bell-shaped velocity profiles. The end-effector orientation sampling is calculated using a computer graphics method slerp for constant angular velocities, ensuring smooth rotations and the shortest path between initial and final orientations. In summary, HARMONIOUS employs state-of-the-art control problem and reactive obstacle avoidance formulation, but moves beyond the state of the art in providing a unified treatment of three different sensory modalities, the number of DoFs controlled, and incorporation of preferred position motivation and orientation error minimization. Beyond what could fit in Tab. I, HARMONIOUS also features self-collision avoidance and the possibility of setting targets (tasks) for the two robot arms independently or jointly (bimanual task). To our knowledge, HARMONIOUS is unparalleled in the scale of the problem and the number of features it can tackle.
III Materials and Methods
Here we present HARMONIOUS – Human-like reactive motion control and multimodal perception for humanoid robots. An overview is provided in Fig. 2. The core of the solution is the motion controller (red box) and the actual QP formulation (magenta) which is online fed by the target trajectory generator (blue box) and the constraints projected from obstacles (green box).
III-A Experimental setup